A man is 24 years older than his son. Three years ago, he was eight years less than his son's age three years from now. How old is the man?
Found 2 solutions by VFBundy, MathTherapy:
Answer by VFBundy(438)

Man's age = m
Son's age = s

Make two equations:
m = s + 24
m - 3 = (s + 3) - 8

Simplify:
m = s + 24
m - 3 = s - 5

Move variables to the left and numbers to the right:
m - s = 24
m - s = - 2

Since both equations say m - s equal different amounts, there is no solution to this problem.

Answer by MathTherapy(10552)

A man is 24 years older than his son. Three years ago, he was eight years less than his son's age three years from now. How old is the man?
Let ma's age be M and son's, S
Then we get: S = M - 24 ------- eq (i)
Also, M - 3 = S + 3 - 8______M - 3 = S - 5______S = M + 2 ------ eq (ii)
MAKES NO SENSE at all!! The son CANNOT be 24 years younger than his father, and at the same time, be 2 years older than his father.
If this was copied correctly, who the H..L writes these problems?
Even 5th, and in some cases, 4th, 3rd, 2nd, or 1st graders are smarter than these supposed math "story-problem" writers.